OpenStudy (anonymous):

use trigonometric form and de moivre's theorem to simplify the expression. (2+2i)^3

OpenStudy (anonymous):



OpenStudy (anonymous):

2+2i= \[\sqrt{8}(\cos(45) +i*\sin(45))\]

OpenStudy (anonymous):

\[(\sqrt{8}(\cos(45)+i*\sin(45))^3 = \sqrt{8}^3*(\cos(45)+i*\sin(45))^3\] This is where good ol' De'Moivre comes in.

OpenStudy (anonymous):

\[=\sqrt{8}^3(\cos(3*45)+i*\sin(3*45))\]

OpenStudy (anonymous):

cos(8*pi/4)? I believe the angle is simple pi/4.

OpenStudy (anonymous):

And then after De'Moivre's, you should be left with cos(3pi/4) + i*sin(3pi/4)

OpenStudy (mertsj):

\[(2+2i) 3 =(8 √ ) 3 (\cos3×pi 4 +isin3×pi 4 )=162 √ (\cos3pi/4+isin3pi/4)=\]

OpenStudy (anonymous):

Yes, that.

OpenStudy (anonymous):

i think i get it, thanks smoothmath!